# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ 3. Methods
In this section, we describe our methods and study procedure. We conducted a task-based semi-structured interview study with 43 participants recruited from civil society organizations, a job agency, and the author’s extended network (Section 3.5), structured into 8 focus groups (3–5 participants) and 12 single interviews. Participants were presented with the study’s employment prediction use case (Section 3.1) and received a collection of explanations about this system (Section 3.2) before solving four tasks and making decisions about the system’s deployment (Section 3.3). The study closed with an interview, lasting 90–120 minutes for focus groups and 60 minutes for single interviews. We analyzed the explanations’ effect, participants’ prioritization of information, and the individual and collective interaction with the explanations (Section 3.4). The university’s research ethics committee approved this study.
3.1. Use case: The employment prediction algorithm
The employment prediction algorithm444More detailed information on the use case has been omitted to adhere to the anonymization policy but will be re-inserted for the final version. is a system meant to support counselors of a public employment agency in assessing job-seekers’ employability. It uses a logistic regression model trained on historical data to assess how demographic features (e.g., age, education, nationality) influence employment chances, producing short-term and long-term employment scores as recommendations. Counselors use these recommendations to decide about adequate support measures for the job-seeker (courses, stabilizing measures, etc.) but can override the system’s recommendations with justification. The system is comparable to algorithmic profiling tools that are used in various countries to assist in job-seeker assessment and resource allocation (Scott et al., 2022; Lopez, 2019; Niklas et al., 2015) and which repeatedly highlight the limits of automation in interpersonal interaction (Bundesagentur für Arbeit, 2021; Niklas et al., 2015; Desiere and Struyven, 2021; Scott et al., 2022). Due to its documentation and exemplification of typical ADM systems’ challenges in public employment, the employment prediction algorithm was thus chosen as the use case for this study.
@@ 3.2. Explanation design
Our explanation design was composed of 36 question-answer pairs about the employment prediction algorithm. Each question belonged to one of four categories, data (format, content, limitations), system details (features, model, examples), usage (operation by and interaction with users), and context (intention of deployment, target group, responsible actors), which were informed by work about the information needs of AI novices (Schmude et al., 2024). Each category was further divided into subtopics with three levels of increasing detail (base level, level 2, level 3). Every explanation contained a question (e.g., ”Who operates the system?”) answered with a brief text or image printed on an A5 paper sheet (Figure 2). Participants first received an explanation overview (Figure 1) and the four base explanations and could request levels 2 and 3 at any time during the explanation phase (as depicted in Figure 3).
The structure of the explanations was intended to i) facilitate uptake by providing information in a modular format and ii) provide different levels of both soundness (information levels) and completeness (subtopics) of information (Kulesza et al., 2013, 2015; Chatti et al., 2022). The selection of subtopics and information levels was guided by work on ADM systems in public employment (Scott et al., 2022; Niklas et al., 2015; Allhutter et al., 2020) and research on the prioritization of information by AI novices (Schmude et al., 2024; Weitz et al., 2024). The self-directed provision of information allowed participants to select explanations according to their interests (Conati et al., 2021) and information needs. The question-answer style was motivated by work on question-driven explanations (Liao et al., 2020) and was intended to encourage engagement. The explanations made use of different techniques that include feature importance, local and global explanations, examples, counterfactuals, and argumentative approaches (Figure 2). Lastly, the explanations were presented in an analog paper format to encourage interactions such as handing around paper sheets, pointing, and reading aloud.
3.3. Study setup and procedure
The study procedure consisted of several steps that are depicted in Figure 3. This section describes the introduction of the use case, the explanation phase comprising exploration, tasks, and the group decision, and the mechanisms for participants’ self-reports on understanding, confidence, and other aspects.
Use case introduction. Participants received initial information about the employment prediction algorithm in the form of a mock newspaper article. The article provided key information and featured the perspectives of employers and employee associations.
@@ 3.3. Study setup and procedure
Exploration phase. After receiving the explanation overview and base level explanations of each category, participants had 15 minutes (focus groups) or 20 minutes (interviews) to become familiar with the structure and mark relevant explanations. Participants could freely decide which explanations to read and request and were encouraged to discuss them.
@@
Tasks. Participants then received a fictional job-seeker case example describing Mr. Harald G.555Details on the inspiration for this case example where omitted to adhere to anonymization policy, but will be re-inserted for the final version. , a fictional job-seeker with a brief backstory and a list of features that would be used to calculate his employment chances. Participants solved four tasks formulated as questions covering i) the job-seeker’s ability to change information stored about him (such as education, physical impairment, etc.), ii) his employability scoring by the employment prediction algorithm, iii) the support measures he would receive, and iv) his ability to contest the decision. All questions had multiple answers with one correct choice and could be answered with information present in the explanations in different categories and detail levels. Whereas tasks 1, 3, and 4 required mostly information retrieval, task 2 could be solved in two ways: by either giving an estimate based on the rough weightings in the system details base explanations or by calculating the precise employment score. Participants had 15 (focus groups) or 20 (single interviews) minutes to solve the tasks. During that time, they could access and request all explanations and discuss possible solutions.
Decision task. Lastly, participants in focus groups had 10 minutes to discuss the system’s deployment and make a collective yes / no decision. If no consensual decision was reached in time, participants either decided themselves how to resolve the situation (e.g., majority vote) or received the option of stating conditions for the system’s deployment.
@@
Self-reports. Participants reported their understanding, decision confidence, and inclusion and voice in the group with 5-point Likert scales, their decision on the system’s deployment with yes / no voting ballots, and their choice of the explanation categories most ”helpful for their understanding” and ”influential to their decisions” by selecting the categories that applied. Self-reports were conducted using a color-coded system to ensure anonymity during focus groups. Each participant was assigned a color and given matching reporting material. During the second self-report, the investigator also asked interview questions about participants’ interaction with the explanations, their understanding processes, the most relevant information, and any additional situational questions. In single interviews, the collective decision and questions about inclusion and voice in the group were omitted.
@@ 3.4. Analysis
All focus groups and interviews were audio-recorded and transcribed. These transcripts, participants’ self-reports during the studies, and the investigators’ field notes, provided the data basis for analysis.
For RQ1-Explanations, we focus on how the explanations helped and impacted participants’ understanding of the ADM system and examine the differences between individual and group settings. For this, we compare participants’ self-reported understanding before and after the explanation phases, their performance in the study tasks, and which and how many explanations were requested. Further, we conduct an inductive thematic analysis (Braun and Clarke, 2006) of participants’ responses to the investigator’s interview questions, capturing articulations about the strengths and weaknesses of the explanations’ design, the impact of the time limit, understanding barriers, and information perceived as most relevant. While the quantitative items thus serve to characterize the diversity of participants’ perceptions (Maxwell, 2010) and facilitate qualitative exploration (Weiss, 1995), they are not intended to invoke ”inference […] of greater generality” (Maxwell, 2010) nor impose a mental model based on variance theory (Patton, 1990).
For RQ2-Deliberation, we first focus on how participants shared and combined information to solve the study tasks and, second, on how explanations enabled group deliberation. For the first part, we conduct an inductive and deductive thematic analysis of the groups’ interaction processes with explanations, connecting our descriptions of these interaction processes to mechanisms for collaborative success and failures as described by Nokes-Malach et al. (2015). For the second part, we compare self-reported decision confidence before and after participants received and discussed the explanations, summarize their feedback about the group setting, and conduct an inductive thematic analysis of focus group conversations in which explanations drove group deliberation and convergence, facilitating the opinion-forming of participants.
@@ 3.5. Participants
Recruitment. Table 1+2 provides an overview of the study participants. Participants were recruited through cooperation with civil society organizations, an employment agency, and the authors’ extended network. Recruitment criteria included full legal age and no technical knowledge about AI systems, i.e., AI novices. All studies were conducted in person in office or public spaces. Participants were compensated with 30€ for participation in focus groups (90–120 minutes) and 20€ for single interviews (60 minutes). Our approach for organizing, composing, and moderating focus groups was informed by Krueger (2004). Concerning the participant sample size, we are guided by research on qualitative methods, which suggests that the number of participants should be determined by code and meaning saturation (Hennink et al., 2017).
Choice of participants. Participants were selected to be representatives of one of three roles: domain experts, decision subjects, or members of the general public. We define domain experts as people who are competent in the field that the ADM system is used in, such as job counselors or advisors (groups 2, 3, 7, and 8). We define decision subjects as people who would potentially be impacted by the system’s decision, such as job-seekers and people who had previously been job-seeking (groups 5 and 6). All remaining participants are considered as members of the general public and were included to test changes in explanation effects and participants’ perceptions (groups 1 and 4). The study was conducted with separate participants in three pilot groups to test and refine the explanation design and study procedure.
@@
| B3 | 46 | University | Social counselor | G2 | 49 | University | DSGVO officer | ||
@@ 4. Results
In this section, we present our results, structured according to our research questions: how question-driven, modular explanations support understanding in individual and group settings (RQ1, Section 4.1) and how groups used explanations666We describe every question and answer pair as one ”explanation” and their entirety as a ”collection of explanations” (Section 3). to solve the study tasks and form opinions about the system’s deployment (RQ2, Section 4.2). Participant labels denote the study setting (focus group: A–H / single interviews: S) and the individual participant, as listed in Tables 1+2.

### v2

@@ 3. Methods
In this section, we describe our methods and study procedure. We conducted a task-based semi-structured interview study with 43 participants (Section 3.6), structured into 8 focus groups with 3–5 participants each and 12 single interviews (Section 3.3). Participants were presented with the study’s employment prediction use case (Section 3.1) and a collection of explanations about this system (Section 3.2) before solving four tasks and deciding about the system’s deployment (Section 3.4). The study closed with an interview, lasting 90–120 minutes for focus groups and 60 minutes for single interviews. We analyzed individual and collective interactions with the explanations, self-reports, and deliberation processes (Section 3.5). The university’s research ethics committee approved this study.
3.1. Use case: The AMS employment prediction algorithm
3.1.1. Description
The AMS algorithm 555AMS stands for the Public Employment Agency (Arbeitsmarktservice). is a system developed to calculate the employability of job-seekers in Austria. It was created by a private company for the Austrian Public Employment Agency between 2015 and 2021 but was never used as a live system and put on hold in 2021 due to privacy objections (Allhutter et al., 2020). The system uses a logistic regression model trained on historical data to predict job-seekers’ employment chances based on demographic features (such as age, education, nationality, etc.) and work history. The outputs are a short-term and long-term employment score for each job-seeker (Gamper et al., 2020). These scores would serve as recommendations for the job-seekers’ counselors at the employment agency to assist in deciding about suitable support measures. Counselors could overwrite the system’s predicted scores of job-seekers but would need to give a reason for doing so (Allhutter et al., 2020; Holl et al., 2018). More information is provided in the appendix.
3.1.2. Choice of use case
Algorithmic tools that assist in assessing job-seekers and resource allocation have been deployed in various countries, including Germany (Bundesagentur für Arbeit, 2021), Austria (Allhutter et al., 2020), Poland (Niklas et al., 2015), and the Netherlands (Desiere and Struyven, 2021). However, the introductions of these applications also repeatedly led to sociotechnical conflicts (Scott et al., 2022). The deployment of the AMS algorithm was motivated by three overarching goals: a) increasing consultation efficiency, b) increasing support measure effectiveness, and c) reducing arbitrariness (Gamper et al., 2020). Detailed reports warned that counselors might over-rely on the algorithm or hesitate to overrule its suggestions (Allhutter et al., 2020). Further, the algorithm’s model and underlying data structure were predicted to discriminate against marginalized groups, who would lack the option to contest the system itself (Lopez, 2019). Transparency and ongoing scrutiny of the algorithm were listed as necessary measures to prevent these risks (Allhutter et al., 2020). As the AMS algorithm represents a larger class of algorithmic decision-making systems that spark public debate around their deployment in public institutions (Raji et al., 2022), it exemplifies how AI systems become matters of public interest and presents a suitable use case for our study.
@@ 3.2.1. Description
The explanation design comprised 36 question-answer pairs about the AMS algorithm. Each question belonged to one of four categories, data (format, content, limitations), system details (features, model, examples), usage (operation by and interaction with users), and context (intention of deployment, target group, responsible actors). Each category was further divided into topics with three levels of increasing detail (base level, level 2, level 3). Every explanation was printed on an A5 paper sheet and contained a question (e.g., ”Who operates the system?”) answered with a brief text or image (cf. Figure 2). Participants first received an explanation overview (Figure 1) and the four base explanations and could request levels 2 and 3 at any time during the explanation phase (as depicted in Figure 3). The explanations were presented in an analog paper format that allowed participants to interact with them physically and that facilitated social interactions, such as sorting, exchanging, pointing, and reading to each other. The full collection of explanations is depicted in the appendix.
3.2.2. Design foundations
The explanation design was intended to allow the users to explore information in a flexible and self-directed manner. To this end, the design used a modular structure, meaning that the explanations were divided into four information categories, which were again subdivided into topics and three levels of detail (base level, level 1, level 2). The four information categories, data, system details, usage, and context were based on research on AI novices’ information needs and covered both technical and sociotechnical system aspects (Schmude et al., 2024). The subdivision of explanation categories into topics and levels of detail organized this broad supply of information while accommodating different needs of information completeness and soundness (Kulesza et al., 2013) and introducing a degree of personalization (Chatti et al., 2022). The goal of the modular design was thus to create explanations that offered information on every aspect of the AI system, from which participants could select the most relevant according to their information needs and preferences. It further aimed to avoid limitations of “groupware” systems (Mandviwalla and Olfman, 1994) by supporting both individual and collaborative interaction, providing multiple user perspectives (e.g., user, decision subject), and synchronizing interaction with the material.
The question-answer style was motivated by explanation design research (Liao et al., 2020, 2021) and was intended to improve user engagement and understanding by matching their thought processes. For example, when users learn that the system uses features to calculate job-seekers’ employment scores, they might ask what the exact weights of these features are and how they are calculated. This corresponds to the three levels of detail in topic A of system details (Figure 1). The explanations further used different explanation methods (Speith, 2022) such as feature importance, local and global explanations, examples, counterfactuals, and argumentative approaches (Figure 2). Information was presented in different formats but mostly relied on textual information and used highlighting, colors, and illustrations to emphasize key points.
3.3. Study procedure
We describe the procedure in the focus group and single interview setting (depicted in Figure 3). Like previous work in XAI (Chiang et al., 2023), we conducted individual and group settings to compare how the social setting would affect participants’ understanding and deliberation processes.
3.3.1. Focus group procedure.
Throughout the study, participants sat together with the investigator and could freely interact with each other. They first completed consent forms and questionnaires about demographics and knowledge about employment (domain knowledge) and AI systems (technical literacy). A round of introductions followed, where each group member stated their name and last interaction with AI to break the ice. The investigator then explained the study procedure and distributed a mock newspaper article introducing the AI use case (included in the appendix). Participants indicated their understanding, deployment decision, and decision confidence for the first time (3.4). The study’s explanation phase followed (including orientation, task, and decision phase), throughout which participants received and kept access to all explanations. In the beginning, the group received an overview of the explanations and all base level explanations, all other explanations could be requested at any time. After 15 minutes, participants received task sheets and had another 15 minutes to complete them, deciding independently whether they wanted to work together or individually. Finally, the group had 10 minutes to make a joint deployment decision (yes/no, with conditions allowed). A second round of individual reports followed, described in the next section, and then the investigator concluded the study. Focus group studies took around 90 minutes.
3.3.2. Single interview procedure.
The majority of the study procedure remained the same in single interviews. However, the explanation phase did not include a group decision phase; instead, the orientation and task phases were prolonged to 20 minutes each to provide individuals with the same total time as focus groups to interact with the explanations and individual reports about the group setting were omitted. The individual study setting was included to analyze how understanding processes changed depending on whether participants worked in a group or alone and whether the explanation design would support both settings. Single interviews took around 60 minutes.
3.4. Study elements
This section provides descriptions and motivation for the study elements: the introduction of the use case, the explanation phase comprising orientation, tasks, and the group decision, and participants’ individual reports.
3.4.1. Use case introduction.
Participants received initial information about the AMS algorithm in the form of a mock newspaper article inspired by an Austrian newspaper publication from 2019 (Szigetvari, 2018). The article provided key information about the system’s basic workings, goals, and deployment context and featured the opinions of employers and employee associations about its merits and risks. The presentation format was chosen to provide an introductory summary of the AI system using a familiar layout and non-technical language while highlighting both the pros and cons of the system’s deployment. Thus, The article served as a basic introduction to the use case, which aimed to approximate the amount of information participants might receive through the media. This way, participants received a baseline of information with which to assess their initial understanding, deployment decision, and decision confidence. Further, this introduction served to outline relevant aspects that could be explored using the explanations.
3.4.2. Explanation phase (orientation, tasks, and group decision).
@@ 3.4. Study elements
For orientation, participants received the explanation overview (Figure 1) and base-level explanations (the first row of explanations in each category) and had 15 minutes (focus groups) or 20 minutes (single interviews) to become familiar with the structure and explore topics of interest. Explanations were provided as single A5 sheets to promote the physical sharing and exchanging of explanations. Participants could freely decide which explanations to request and read, and whether to share and discuss information with others.
@@
For the tasks, participants received the case of Mr. Harald G.666The case example was inspired by Allhutter et al. (2020) and adapted to this study, as depicted in the appendix., a fictional job-seeker with a brief backstory and a list of features. Participants had 15 minutes (focus groups) or 20 minutes (single interviews) to solve four tasks pertaining to this case. All questions could be answered with information from explanations in different categories and levels of detail. Whereas tasks 1, 3, and 4 required locating information, task 2 could be solved in two ways (aside from guessing): by either giving an estimate based on the rough weightings in the system details base explanations or calculating the precise employment score. Participants could access and request all explanations and discuss possible solutions. These were the four tasks (correct answers underlined):
Task 1: Can Harald change the data stored about him (e.g., to correct it)? (yes / no)
Task 2: In which group of employment chance does the system categorize Harald?
(high (¿66%) / medium (¡66% & ¿25%) / low (¡25%))
Task 3: Which support measures will Harald receive?
(qualifying measures, such as courses and training / stabilization and increased supervision / none)
Task 4: Can Harald appeal against this decision? (yes / no)
For the group decision, focus group participants had 10 minutes to discuss the system’s deployment and were asked to collectively decide whether to accept or reject it. This was meant to simulate a small referendum in which each participant’s voice counted for the final outcome. If no consensual decision was reached in time, participants were asked how the situation should be resolved (e.g., majority vote). They were further made aware of the option to state conditions for the system’s deployment.
3.4.3. Individual reports I and II
@@
Participants were asked for individual reports before and after the explanation phase. At both points, participants reported understanding (5-point scale), deployment decision (yes / no), and decision confidence (5-point scale) to examine the effect of the explanation phase. In report II, participants also reported the explanation categories that were most helpful for their understanding (multiple choice) and most influential to their decisions (multiple choice); focus group participants additionally reported perceived inclusion in the group and the discussions’ influence on their decision (5-point scales). During report II, the investigator asked participants interview questions about their interaction with the explanations, understanding processes, prioritized information, and situational aspects. The list of interview questions is included in the appendix.
To prevent influence between participants’ reports, individual reports in focus groups were conducted anonymously and re-assigned by the study examiner using a color-coded reporting system (Figure 4). Each participant was assigned a color and received the material for all ten individual reports. Participants took the corresponding paper slip for each report, wrote their answer, and threw it in a gathering container that hid it from view. The gathered reports were then collected and recorded by the investigator at the end of the study.
@@ 3.5. Analysis
All focus groups and interviews were audio-recorded and transcribed. These transcripts provided the data basis for the thematic analysis. Participants’ individual reports, task solutions, decisions, and the investigators’ field notes provided further data for within- and between-subject comparisons of understanding and decision-making.
3.5.1. Thematic analysis.
For both research questions, we conducted thematic analysis (Braun and Clarke, 2006) of participants’ articulations to develop a qualitative account of their understanding and decision-making processes. To this end, we developed inductive codes in the first pass and refined them in the second pass on the transcriptions. The resulting inductive code base was structured along the overarching categories of understanding, deliberation (decision-making processes, arguments), opinions (e.g., about AI and policy choices), and experiences (e.g., anecdotes and lived situations). The full code-book is provided in the appendix. We further highlight that while the quantitative items in participants’ self-reports serve to characterize the diversity of participants’ perceptions and facilitate qualitative exploration (Weiss, 1995), they are not intended to invoke “inference […] of greater generality” (Maxwell, 2010) nor impose a mental model based on variance theory (Patton, 1990).
3.5.2. RQ1-Explanations.
RQ1 focuses on the explanations’ impact on participants’ understanding of the study’s use case and the differences between individual and group settings. We employ a triangulation approach (Carter et al., 2014) by using three ways of analyzing the explanations’ effect on participant understanding. Firstly, comparing participants’ individual reports before and after the explanation phase was a subjective indicator of changes in their understanding and decision confidence. Secondly, participants’ answers to the four study tasks indicated their factual or testable understanding. Lastly, participants’ verbal reports during and after the explanation phase were used to analyze their understanding processes and barriers thematically. In a deductive analysis, we further compared their interactions to mechanisms of “collaborative success and failure” (Nokes-Malach et al., 2015) and the “six facets of understanding” (Wiggins and McTighe, 2005). With this three-part combination, we examined participants’ subjective understanding, their information gain, and the cognitive processes of their understanding. This choice was motivated by educational psychology research outlining that understanding cannot only be elicited through test questionnaires (Sato et al., 2019) but involves emotional (Wiggins and McTighe, 2005) and meta-cognitive processes (Veenman et al., 2006) that are equally important. Our focus on understanding is motivated by previous XAI research, which highlights the importance of understanding in decision-making processes (Langer et al., 2021; Shin, 2023; Hoffman et al., 2023).
3.5.3. RQ2-Deliberation.
RQ2 focuses on how participants formed opinions about the AMS algorithm, weighed the pros and cons of its deployment, and settled on a deployment decision. To this end, we compare participants’ decision confidence before and after the explanation phase. We further conduct an inductive and deductive thematic analysis of participants’ interactions in both settings to connect them to the “elements of deliberation” (Stromer-Galley, 2007) – a set of characteristics that outline deliberation processes. Based on this, we analyze when participants used arguments (grounded, defensible positions), opinions (personal judgments on things, values, states), and personal experiences (Stromer-Galley, 2007; Mercier and Landemore, 2012) to consider the system’s deployment. For single interviews, we examine participants’ responses to interview questions during the study to examine their reasoning process and “internal deliberation” (Mercier and Landemore, 2012). Lastly, to account for one of the most prevalent cognitive biases in group settings, we examine focus groups for occurrences of groupthink (Janis, 1971; Baron, 2005) – an effect that sets in when concurrence-seeking in groups overrides realistic argumentation and discussion.
@@ 3.6. Participants
3.6.1. Recruitment.
Table 1 and 2 provide an overview of the study participants. Participants were recruited through cooperation with civil society organizations, an employment agency, public calls for participation, and the authors’ extended network. For focus groups, the authors contacted staff from these organizations known from previous studies or used channels of general inquiry to describe the study and invite participation. Interested organizations all offered to support the recruitment process by coordinating with the authors on selecting and inviting potential participants and finding a place and time to conduct the studies. Groups A, B, C, E, F, G, and H were organized this way. Group D was recruited through the authors’ network and was equally composed of people who had previously been job-seeking. For individual studies, participants were recruited using the same channels, and calls for participation were further advertised on screens and information boards throughout different city districts. All studies were conducted in person in office or public spaces. Participants were compensated with 30€ for participation in focus groups (90–120 minutes) and 20€ for single interviews (60 minutes). Our approach for organizing, composing, and moderating focus groups was informed by Krueger (2004). Concerning the participant sample size, we are guided by research on qualitative methods, which suggests that the number of participants should be determined by code and meaning saturation (Hennink et al., 2017).
3.6.2. Recruitment criteria.
Participants were required to be of full legal age and AI novices, i.e., to have no technical knowledge or expertise about machine learning systems as described in Section 2.2. These criteria were screened in a pre-questionnaire before invitation to the study using two questions: “How would you rate your knowledge of algorithms?” and “How would you rate your knowledge of artificial intelligence (AI)?”. Each question could be answered on a scale corresponding to “no knowledge at all” to “professional and detailed knowledge”. Here, the first question elicited technical knowledge, as participants familiar with AI tools might have rated their AI expertise as high but were unlikely to know about algorithms without a strong technical interest or background.
3.6.3. Group composition.
Participants were further selected to be representatives of one of three roles: domain experts, decision subjects, or members of the general public. We define domain experts as people who are competent in the field that the AI system is used in, such as job counselors or advisors (groups B, C, G, and H). We define decision subjects as people who would potentially be impacted by the system’s decision, such as job-seekers and people who had previously been job-seeking (groups E and F). All remaining participants are considered members of the general public and were included to test changes in explanation effects and participants’ perceptions (groups A and D). The study was conducted with separate participants in three pilot groups to test and refine the explanation design and study procedure.
@@
B3 | 46 | University | Social counselor | G2 | 49 | University | GDPR officer | ||
@@ 4. Results
In this section, we present our results as answers to our research questions: How question-driven, modular explanations777We note again that with “explanation” we mean a question and answer pair and with “explanations” we mean the collection of all 36 explanations (Section 3). support understanding in individual and group settings (RQ1, Section 4.1) and how AI novices used explanations to form opinions and decide about the system’s deployment (RQ2, Section 4.2). Participant labels denote the study setting (focus group: A–H / single interviews: S) and the participant ID, as listed in Tables 1+2. To distinguish themes in the analysis, inductive themes are italicized, while ‘deductive themes’ are put in quotes.